• Show that the tension T of the string is related to the fundamental frequency f1 by where L is the length of the string, and u is the linear mass of the string.
Answer #1

here

the fundamental frequency is

f1 = v / 2*L

and the velocity v= sqrt(T / (m/L))

then put the value of v

f1 = sqrt(T/(m/L)) / 2*L